a) If the interest rate in the United Kingdom is 8%, the interest rate in the U.S. is 10%, and the spot exchange rate is 1.35 dollar per pound. If interest rate parity holds, what is the expected future exchange rate?

b) Alternatively, using the same interest rates as above, suppose the expected future exchange rate is 1.35 dollar per pound. What is the spot exchange rate?
Answer #1

a) Using interest parity condition

RUS = RUK = (Ee - E)/E

10% = 8% + (Ee - 1.35)/1.35

0.02 = (Ee - 1.35)/1.35

Ee = 0.02*1.35 + 1.35 = 1.377

Hence,  expected future exchange rate is 1.377 dollars per pound

b) Using interest parity condition

RUS = RUK = (Ee - E)/E

10% = 8% + (1.35 - E)/E

0.02 = (1.35 - E)/E

0.02E =1.35E - E^2

E^2 - 1.33E = 0

E = 1.33

Hence, spot exchange rate is 1.33 dollars per pound